OpenStudy (anonymous):

need some one to check and make sure i did number 7 correctly

OpenStudy (anonymous):

srry bout it being sideways

OpenStudy (anonymous):

B is true in number 7

OpenStudy (anonymous):

C is the only one not true

OpenStudy (anonymous):

soo im right???

OpenStudy (anonymous):

no, it asked which is *not* valid. B is valid which is what you circled. C is the only statement that is not valid

OpenStudy (anonymous):

oh its asking which one is wrong ok ok

OpenStudy (anonymous):

yep A,B,D are true

OpenStudy (anonymous):

6. is B

OpenStudy (anonymous):

yea (2-n)*180=3*180=540

OpenStudy (anonymous):

can i get sum help on # 9 now u guys got any idea on dat one???

OpenStudy (anonymous):

SAS

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

C. SAS

OpenStudy (anonymous):

Because C is a vertical angle. They are congruent on both sides.

OpenStudy (anonymous):

C (the angle in the diagram)

OpenStudy (anonymous):

And you know the other 2 sides are congruent so it would be Side-Angle-Side in congruence.

OpenStudy (anonymous):

ok ok how bout # 5

OpenStudy (anonymous):

5 inches

OpenStudy (anonymous):

i think 8 is wrong the converse should be D.

OpenStudy (anonymous):

lucas can you look at 8 and tell me if you agree

OpenStudy (anonymous):

statement: if p then q converse: if q then p

OpenStudy (anonymous):

very hard to see if u can rota

OpenStudy (anonymous):

Yeah it's D. Haha hurting my neck trying to help with some math....

OpenStudy (anonymous):

umm how bout # 5

OpenStudy (anonymous):

# 5 is B

OpenStudy (anonymous):

why wat did i do wrong

OpenStudy (anonymous):

Think about it, if it's a circumcenter that means each line is bisected. So line KW is congruent to line WL which is the other half of the full line, line KL.